The velocity of an H-atom in ground state is 2.2*10^6 m/s.The de-broigle wavelength of the electron would be???

Pls provide the exact answer someone...

Grade:12

1 Answers

vikas askiitian expert
13 years ago

lamda = h/mv

h = 6.6*10-34

m = 9.1*10-31

v = 2.2*106 m/s

lamda = 3.29*10-10m  = 3.3 angstromes unit